# Efficient Computation of Whole-Body Control Utilizing Simplified Whole-Body Dynamics via Centroidal Dynamics

### v1

@@ 4 End effector position & orientation $6\times k\text{ eq}$
TABLE II: Task Configuration with Proposed Method
@@ 2 End effector position & orientation $6\times k\text{ eq}$
V-B Result with Proposed Method
The task for comparative experiments is configured as seen in Table I. These tasks were constructed similarly to the tasks in [22], for bipedal robot walking. Table I shows the various task hierarchies of LQP formulation configured using the original whole-body model. Since the walking is consists of single contact and double contact scenarios, Table II shows how the tasks from Table I are divided into two parts utilizing the proposed method. The result measured the computation time in single support ($c=1$, $k=2$) and double support ($c=2$, $k=1$) scenarios occurring during walking. The tasks configured in this manner were structured in an LQP formulation and computed through the sequential method.
The results of comparing the computation times for solving the LQP problem using the sequential method can be observed in Table LABEL:tableComputenew.
Table LABEL:tableComputenew compares the average computation times of the conventional method and the proposed method during double contact and single contact scenarios. In both scenarios, the proposed method demonstrated an improvement in computational performance, with a 299% increase in the single support scenario and a 222% increase in the double support scenario, indicating enhanced performance compared to the original computation times.
Figure 4: Tracking errors of COM and end-effector positions using conventional and proposed methods. The tracking error was measured by controlling the robot during double support.
The trajectory tracking errors resulting from task control using the proposed method and the conventional method were measured and compared. The COM and end-effector of the arm are controlled at the same time. The control errors can be observed in Figure 4. Similar magnitudes of control error were observed in both methods. Task space control performance varies with the tuning of the weighting matrix, making it challenging to compare the two methods under perfectly identical conditions. Nevertheless, it can be seen that both methods achieve a comparable level of control.
TABLE III: Comparison of Computation Times with Support types
@@ V-B Result with Proposed Method
V-C Compuational Performance Analysis by DOF
The proposed method was also analyzed to determine its impact on the computational performance of LQP based on the robot’s DOF. With both feet in contact, the benchmark examined how changing the DOF in the upper body affected computational performance.
Task configuration for the comparison is also set up similar to Table I and Table II. The difference lies in the task within the unconstrained chain. The end-effector of the arm is defined as the task space.
Benchmark tests were conducted by varying the robot model’s DOF from 20 to 45, and the results are displayed in Figure 5. When tested on a 20-DOF robot model, the proposed method took 571 $\mu$s, while the conventional method required an average computation time of 754 $\mu$s, showing a 132% improvement in computational performance. This computational performance improvement increases with the robot model’s DOF. When tested on a 45-DOF robot model, the proposed method took only 1319 $\mu$s, while the conventional method required 4081 $\mu$s, resulting in a 310% improvement in computational efficiency.
@@ V-C Compuational Performance Analysis by DOF
VI CONCLUSIONS
In this study, a novel method for constructing reduced-dimension rigid body dynamics is proposed. This method employs a reduced-dimension dynamics based on constraints of the kinematic chain, partially projecting the unconstrained kinematic chain into it’s centroidal space, aiming to improve computational efficiency. The proposed method offers a unique approach by dividing the model into two parts with lower DOF, thereby enhancing computational efficiency by handling lower-DOF models in each sequence of the problem.
@@ VI CONCLUSIONS
When the proposed method was applied to the LQP formulation, which is commonly used in other studies, and computed using the sequential method, a 222 % improvement in compuational efficiency was observed for the 33-DOF model at double support, and a 310% improvement was observed for the 45-DOF model. This demonstrates a significant performance improvement with increasing DOF. The control performance was also found to be similar between the two methods, demonstrating that computational time can be reduced without compromising control performance.

### v2

@@ 4 End effector position & orientation $6\times k\text{ eq}$
Table 2: Task Configuration with Proposed Method
@@ 2 End effector position & orientation $6\times k\text{ eq}$
5.2 Result with Proposed Method
The task for comparative experiments is configured as seen in Table 1. These tasks were constructed similarly to the tasks in [18], for bipedal robot walking. Table 1 shows the various task hierarchies of LQP formulation configured using the original whole-body model. Since the walking is consists of single contact and double contact scenarios, Table 2 shows how the tasks from Table 1 are divided into two parts utilizing the proposed method. The result measured the computation time in single support ($c=1$, $k=2$) and double support ($c=2$, $k=1$) scenarios occurring during walking. The tasks configured in this manner were structured in an LQP formulation and computed through the sequential method.
Table LABEL:tableComputenew compares the average computation times of the conventional and proposed methods during double and single contact scenarios. The proposed method demonstrated a computation time decrease of 66.5% in the single support scenario and 54.9% in the double support scenario, indicating enhanced computational efficiency.
Figure 4: Tracking errors of COM and end-effector positions using conventional and proposed methods in the double support.
The trajectory tracking errors of the proposed and conventional methods were measured while simultaneously controlling the COM and the arm’s end-effector. As shown in Figure 4, both methods exhibited similar magnitudes of control error. Although task space control performance depends on the tuning of the weighting matrix—making exact comparisons challenging—both methods achieved comparable levels of control.
Table 3: Computation Time Results by Support types
@@ 5.2 Result with Proposed Method
5.3 Compuational Performance Analysis by DOF
The proposed method was analyzed for its impact on the computational efficiency of LQP based on the DOF of the robot’s unconstrained chain. With both feet in contact, benchmarks examined how varying the unconstrained chain’s DOF affected computational efficiency. The task configuration for the comparison is similar to that in Tables 1 and 2, except for the task within the unconstrained chain. Specifically, the arm’s end-effector is defined as the task space.
@@ 5.3 Compuational Performance Analysis by DOF
Figure 5: Average computation time comparison by DOF of the robot, and the relative computation time percentage. The computation time of proposed method is the sum of LQP 1 and LQP 2.
6 CONCLUSIONS
This method employs reduced-dimension dynamics based on the constraints of the kinematic chain by partially projecting the unconstrained kinematic chain into its centroidal space, aiming to improve computational efficiency. The proposed method offers a unique approach by dividing the model into two parts with lower DOF, thereby enhancing computational efficiency by handling lower-DOF models in each sequence of the problem.
@@ 6 CONCLUSIONS
When the proposed method was applied to the LQP formulation, which is commonly used in other studies, and computed using the sequential method, a 54.9 % decrease in computation time was observed for the 33-DOF model at double support, and a 67.7% decrease was observed for the 45-DOF model. This demonstrates a significant efficiency improvement with increasing DOF. The control performance was also found to be similar between the two methods, demonstrating that computational time can be reduced without compromising control performance.
@@
CONFLICT OF INTEREST
The authors declare that there is no competing financial interest or personal relationship that could have appeared to influence the work reported in this paper.